OpenStudy (anonymous):

Solve equation: 1/2lg(3x+1)-1=lg(x-3)-lg(5)

OpenStudy (anonymous):

x>3

OpenStudy (anonymous):

and

OpenStudy (anonymous):

Please show steps

OpenStudy (anonymous):

sorry, this is the first time, so I.... fluent in use tools and my English is very bad

OpenStudy (anonymous):

\[\frac{sqr(3x+1) }{ 2 } = x- 3\]

OpenStudy (anonymous):

well, now could you solve it?

OpenStudy (anonymous):

yes, but maby You can write how to get: sqr(3x+1)/2=x−3

OpenStudy (anonymous):

2lg(3x+1) = lg(3x+1)^2 , ok?

OpenStudy (anonymous):

lg(x-3)-lg(5)= lg[(x-3)/5], ok?

OpenStudy (anonymous):

and lg a = lg b <=> a=b, ok?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

were you clear?

OpenStudy (anonymous):

well, now I must go to bed, bb!